Grade 9Algebra

Can someone explain how to solve this? If a and b are positive integers, and a + b = a/b + b/a, find the value of a^2+b^2?
The answer is 2.

Raghav
9 Years agoGrade 9

1 Answer

Tanishq Chourishi
9 Years ago
as we rearrange the given equation we get (a+b)ab=a2+b2
====>
a2b=b2a=a2+b2
a2(b-1)+b2(a-1)=0  now as its given a and b are positive integers they can also be called as natural numbers...i.e 1,2,3 ,4.....
therefore a-1 willbe greter than equal to 0 
simmilar for b-1...
now let a2(b-1)=x   ;b2(a-1)=y
now our derived equation says that x + y =zero.. where x,y>=0
 
 which is only possible when x=y=0
 thus from a2(b-1) =0  ==>a=0 or b=1
         from b2(a-1)=0  ==>b=0 or a=1
by  taking the common solution from the above satementswe get a=b=1
 
thus a2+b2=1+1=2.......